If Cot θ= 12/5 , and

Cos θ= 1/tan θ. Evaluate;

Cos2θ/sin θ + Cos θ


1
Expert's answer
2021-12-13T14:21:27-0500

"\cot\theta=12/5=>\tan \theta=1/\cot \theta=5/12"

1.

If "\cos θ= 1/\tan θ," then "\cos θ= 1/\tan θ=12/5."

There is no solution, because "-1\leq \cos \theta\leq 1" for "\theta \in \R."


2.

Suppose that the condition "\cos θ= 1/\tan θ" is False.

Then


"1+\cot^2\theta=\dfrac{1}{\sin^2\theta}"

"\sin ^2 \theta=\dfrac{1}{1+\cot^2\theta}=\dfrac{1}{1+(12/5)^2}=\dfrac{25}{169}"

"\cos ^2\theta=1-\sin^2\theta=1-\dfrac{25}{169}=\dfrac{144}{169}"


"\dfrac{\cos(2\theta)}{\cos \theta+\sin \theta}=\dfrac{\cos^2 \theta-\sin^2 \theta}{\cos \theta+\sin \theta}"

"=\dfrac{(\cos \theta-\sin \theta)(\cos \theta+\sin \theta)}{\cos \theta+\sin \theta}"

"=\cos \theta-\sin \theta,\cos \theta\not=-\sin \theta"

i)


"\sin\theta>0, \cos \theta>0"

Then


"\sin \theta=\sqrt{\dfrac{25}{169}}=\dfrac{5}{13}"

"\cos \theta=\sqrt{\dfrac{144}{169}}=\dfrac{12}{13}"

"\dfrac{\cos(2\theta)}{\cos \theta+\sin \theta}=\cos \theta-\sin \theta"

"=\dfrac{12}{13}-\dfrac{5}{13}=\dfrac{8}{13}"

ii)


"\sin\theta<0, \cos \theta<0"

Then


"\sin \theta=-\sqrt{\dfrac{25}{169}}=-\dfrac{5}{13}"

"\cos \theta=-\sqrt{\dfrac{144}{169}}=-\dfrac{12}{13}"

"\dfrac{\cos(2\theta)}{\cos \theta+\sin \theta}=\cos \theta-\sin \theta"

"=-\dfrac{12}{13}-(-\dfrac{5}{13})=-\dfrac{8}{13}"